OpenStudy (anonymous):

Write the equation of a circle with a center at (0, 0) and a radius of 5.

hartnn (hartnn):

the equation of circle with centre as origin and radius = r is \(\large x^2+y^2 = r^2\) here, just plug in r= 5 :)